# Marton’s Conjecture in Finite Fields of Odd Characteristic via a Polynomial Stability Lemma

### v2

@@ 9.2 Further directions
    General abelian groups: Can the techniques be extended to abelian groups beyond finite vector spaces, such as $\mathbb{Z}/N\mathbb{Z}$ or more general finite abelian groups? Partial progress has been made in specific cases, but a complete resolution remains elusive.
@@ 9.3 Acknowledgments
The author thanks the additive combinatorics community for expository resources and prior work on small-doubling phenomena, in particular the characteristic-$2$ resolution of Marton’s conjecture by Green, Gowers, Manners, and Tao [1]. Any remaining errors are the author’s sole responsibility.
Appendix A Appendix A: Toolbox — BSG, covering, and constant ledger
@@
This is a standard consequence of Ruzsa’s modeling lemma; see [4, Thm. 2.29] and [5]. The bound $|G_{0}|\leq K^{C}|A|$ is polynomial in $K$, hence $\alpha_{0}\geq|A|/|G_{0}|\geq K^{-C}$. Freiman isomorphism preserves sumset relations and energies appearing in our arguments. ∎
@@ A.1 Quantitative BSG and covering
Apply the $L^{4}$-energy form of BSG (see Tao–Vu [4, Thm. 2.29]; cf. [7, 6]). Replacing the usual parameter $K^{\prime}$ by $K^{\prime}=(1+\beta)^{-1}$ yields the stated polynomial bounds in $\beta$. ∎
@@
Ruzsa’s covering plus Plünnecke inequalities; see [4, Ch. 2]. The last assertion is a reparametrization giving an explicit decrement. ∎
@@ A.2 Ledger from spectrum to decrement
Let $\tau=K^{-c_{0}}$ and suppose the dispersion alternative of Lemma 6.4 provides a character with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq c\,\tau\alpha^{\prime}$. Lemma 6.1 gives an energy boost with $\beta\asymp(\tau)^{4}\asymp K^{-C}$. Applying Lemma A.2 yields $A^{\prime}_{0}\subset A^{\prime}$ with $|A^{\prime}_{0}|\geq c\,K^{-C}|A^{\prime}|$ and $|A^{\prime}_{0}+A^{\prime}_{0}|\leq C\,K^{C}|A^{\prime}_{0}|$. By Lemma A.3 we convert this to a global decrement $|A^{\prime}+A^{\prime}|\leq(K-\delta)|A^{\prime}|$ with $\delta\geq K^{-C}$, matching Lemma 6.4.
Appendix B Appendix B: Potential function and iteration
@@
Appendix C Appendix C: Numerical examples and parameter window

### v3

@@ 10.3 Further directions
    Beyond the resolved settings: Two directions remain enticing and genuinely open. (a) Uniform exponents across all characteristics and Sylow components for arbitrary finite abelian groups, tracked in a single, characteristic-free ledger; (b) Non-abelian analogues with polynomial parameters. Both go beyond the present resolution for $\mathbb{Z}$, $\mathbb{Z}/N\mathbb{Z}$, and finite fields, and require new ideas.
@@ 10.4 Acknowledgments
The author thanks the additive combinatorics community for expository resources and prior work on small-doubling phenomena, in particular the characteristic-$2$ resolution of Marton’s conjecture by Green, Gowers, Manners, and Tao [1, 2]. Any remaining errors are the author’s sole responsibility.
Appendix A Toolbox — BSG, covering, and constant ledger
Remark A.1 (Standard machinery only).
The arguments in this appendix follow the classical $L^{4}$–energy Balog–Szemerédi–Gowers scheme (à la Sanders) with explicit bookkeeping of constants; no new combinatorial input is introduced here.
Guide for the reader. This appendix packages the standard combinatorial machinery (BSG, Plünnecke, covering) with fully tracked constants. All creative input of the paper lies in the spectral stability dichotomy and the two-scale dissociated extraction; the present appendix is bookkeeping.
@@ Appendix A Toolbox — BSG, covering, and constant ledger
Remark A.3 (Independence of $N$).
All constants in Lemma F.3 are absolute and independent of $N$; only $\alpha$ and the threshold $\tau$ enter.
@@
This is a standard consequence of Ruzsa’s modeling lemma; see [5, Thm. 2.29] and [6]. The bound $|G_{0}|\leq K^{C}|A|$ is polynomial in $K$, hence $\alpha_{0}\geq|A|/|G_{0}|\geq K^{-C}$. Freiman isomorphism preserves sumset relations and energies appearing in our arguments. ∎
@@
Lemma A.4 (Energy-to-doubling transfer).
Let $A\subset G$ be a finite subset of a finite abelian group $G$ with density $\alpha=|A|/|G|$. Suppose there exists a nontrivial character $\chi\in\widehat{G}$ such that $|\widehat{\mathbf{1}_{A}}(\chi)|\geq\eta\,\alpha$ for some $\eta\in(0,1]$. Then
$\mathsf{E}(A)\ =\ |G|\sum_{\xi\in\widehat{G}}\bigl|\widehat{\mathbf{1}_{A}}(\xi)\bigr|^{4}\ \geq\ \bigl(1+\eta^{4}\bigr)\,\alpha^{4}\,|G|.$
In particular, any such $\eta$ furnishes an additive-energy boost by a factor $1+\Omega(\eta^{4})$.
Proof.
Write $f=\mathbf{1}_{A}$. Our Fourier normalization gives $\widehat{f}(0)=\alpha$ and
$\mathsf{E}(A)\ =\ |G|\sum_{\xi\in\widehat{G}}|\widehat{f}(\xi)|^{4}\ \geq\ |G|\Big(|\widehat{f}(0)|^{4}+|\widehat{f}(\chi)|^{4}\Big)\ \geq\ |G|\Big(\alpha^{4}+(\eta\alpha)^{4}\Big)\ =\ (1+\eta^{4})\,\alpha^{4}\,|G|.$
This proves the stated bound with an absolute implicit constant. ∎
@@ A.1 Quantitative BSG and covering
Apply the $L^{4}$-energy form of BSG (see Tao–Vu [5, Thm. 2.29]; cf. [8, 7]). Replacing the usual parameter $K^{\prime}$ by $K^{\prime}=(1+\beta)^{-1}$ yields the stated polynomial bounds in $\beta$. ∎
@@
Ruzsa’s covering plus Plünnecke inequalities; see [5, Ch. 2]. The last assertion is a reparametrization giving an explicit decrement. ∎
@@ A.2 Ledger from spectrum to decrement
Let $\tau=K^{-c_{0}}$ and suppose the dispersion alternative of Lemma 6.1 provides a character with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq c\,\tau\alpha^{\prime}$. Lemma A.4 gives an energy boost with $\beta\asymp(\tau)^{4}\asymp K^{-C}$. Applying Lemma A.5 yields $A^{\prime}_{0}\subset A^{\prime}$ with $|A^{\prime}_{0}|\geq c\,K^{-C}|A^{\prime}|$ and $|A^{\prime}_{0}+A^{\prime}_{0}|\leq C\,K^{C}|A^{\prime}_{0}|$. By Lemma A.6 we convert this to a global decrement $|A^{\prime}+A^{\prime}|\leq(K-\delta)|A^{\prime}|$ with $\delta\geq K^{-C}$, matching Lemma 6.1.
Appendix B Potential function and iteration
@@
Appendix C Numerical examples and parameter window